OpenStudy (anonymous):

solve the equation x^square root logx=10^8

OpenStudy (anonymous):

x^sqrt(log(x)) = 10^8 sqrt(log(x)) * log(x) = 8 sqrt((log(x))^3) = 8 (log(x))^3 = 8^2 = 64 log(x) = 64^(1/3) = 4 x = 10^4

OpenStudy (anonymous):

Thank You!